# Bimanual Regrasp Planning and Control for Eliminating Object Pose Uncertainty

### v1

@@ Abstract
Precisely grasping an object is a challenging task due to pose uncertainties. Conventional methods have used cameras and fixtures to reduce object uncertainty. They are effective but require intensive preparation, such as designing jigs based on the object geometry and calibrating cameras with high-precision tools fabricated using lasers. In this study, we propose a method to reduce the uncertainty of the position and orientation of a grasped object without using a fixture or a camera. Our method is based on the concepts that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact. Three orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three orthogonal grasps of two robotic arms to eliminate uncertainties in the object pose. We evaluated the proposed method on different initial object uncertainties and verified that the method have satisfactory repeatability accuracy. It outperforms an AR marker detection method implemented using cameras and laser jet printers under standard laboratory conditions.
@@
Grasping Uncertainty, Bimanual Manipulation
@@ I Introduction
Asignificant challenge in robotic manipulation lies in addressing the uncertainties associated with object grasping. The uncertainties often arise from errors in environmental registration, inaccuracies in object pose recognition, and unbalanced contact during grasping that leads to pose deviations. The uncertainties can result in discrepancies between the actual and expected pose of objects or tools, potentially causing task failures. Existing methods used to address uncertainty problems typically fall into two categories: Sensor-based correction methods [1][2] and fixture-based constraint methods [3][4]. Sensor-based methods leverage additional hardware, such as force sensors or tactile sensors, to detect and mitigate uncertainties. Fixture-based methods, on the other hand, rely on structuring the environment or using special-purpose structures to constrain object pose. However, both approaches have notable limitations, as they often require additional hardware adjustments or modifications to the environment.
Figure 1: (a) A workpiece is difficult to be accurately grasped without fixture-based positioning. (b) Flat finger pads of parallel grippers can help reduce uncertainty in the hand opening/closing direction. After grasping, the object’s uncertainty is constrained within the gray plane shown in the subfigure. (c) Three orthogonal grasps by parallel grippers constrain an exact pose. This paper proposes a bimanual regrasp planning method that generates regrasping motion sequence by consecutively executing three orthogonal grasps to eliminate object pose and grasp uncertainty.
@@
This paper proposes a novel bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to eliminate uncertainties. The proposed approach does not require external fixtures or additional sensors, making it more flexible for diverse robotic manipulation scenarios. The fundamental concepts of the proposed method are illustrated in Fig. 1(b) and (c). It is based on the observation that the flat finger pads of parallel grippers can reduce uncertainty along the hand’s opening/closing direction through flat surface contact. Three such grasps collectively constrain the object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting mutually orthogonal grasp triples to simultaneously eliminate uncertainties in the object pose. However, grasp planning alone does not fully achieve our goal, as simply executing the planned motion sequence using position control may exert excessive grasping forces on the uncertain object and lead to deformation or even damage. To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors embedded in the robot arm’s wrists to adjust the grasping pose and conform to the uncertain object pose. By integrating regrasp planning with admittance control, our method iteratively eliminates uncertainties along three orthogonal axes, thus achieving an accurate object pose estimation.
In the experiments, we implemented the proposed method and evaluated its performance on different initial object uncertainties. The results demonstrate that the method achieves satisfactory accuracy without the need for external fixtures or additional sensors. Furthermore, we compared the accuracy of the proposed method with that of AR-based detection. The results indicate that our method outperforms the AR marker detection implemented using cameras and laser jet printers under standard laboratory conditions. The proposed approach leverages the robot’s inherent geometric constraints and force sensors to eliminate uncertainties in the grasped object. It improves task success rates by addressing uncertainties in a systematic and hardware-efficient manner.
@@ II Related Work
The primarily related work of this paper focuses on regrasp planning and the elimination of grasping errors. Regrasp planning aims at automatically generating grasp sequences to accomplish object manipulation. It involves two main categories: One aims to identify discrete pick-and-place sequences. For example, Liu et al. [7] developed regrasp planning using multiple mobile robots. The method achieves flipping manipulation with the help of regrasping. Kim et al. [8] developed a randomized planning method for regrasp planning. Baek et al. [9] considered non-prehensile grasping in regrasp planning and secured power grasp for stable holding. Qin et al. [10] presented a regrasp planning method for installing a deformable linear object. Rigidity constraints were considered when planning regrasp motion. Mitrano et al. [11] studied a similar problem and proposed considering avoiding environment-grasp-body loops to accelerate grasp selection and planning. Sundaralingam et al. [12] extended regrasp planning to multi-fingered hand. Murooka et al. [13] proposed the foot-step and regrasping planning method for a humanoid robot to simultaneously determine foot steps and hand switch. The other focuses on continuous adjustments. For instance, Maximo et al. [14] presented the grasp space method for continuous regrasp (rolling or sliding) that maintains force closures. Hu et al. [15] formulated regrasp planning as a multi-modal problem, where three manifolds, including the regrasp manifold, the transfer manifold, and the slide manifold, were planned across to determine action sequences and thus plan motion. Cheng et al. [16] further proposed the contact mode method to guide manipulation planning across multiple manifolds. The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it should be categorized under the first type of regrasp planning.
@@
One of the most critical challenges in regrasp planning is the planning speed. Researchers have explored various methods to enhance efficiency. For example, Cho et al. [17] developed a fast regrasp planning algorithm using look-up tables and two-finger parallel grippers. Chiu et al. [18] proposed a reinforcement learning method to accelerate the planning of regrasp poses. Their actions were defined in the local frame of a robot hand, thereby resolved the dependencies on robotic kinematics. Wada et al. [19] significantly improved planning speed by using a neural network to filter waypoint poses and quickly select grasp poses. Xu et al. [20] proposed learning reorientation object poses and then plan regrasp motion while considering the learned middle reorientation poses. Learning-based methods have demonstrated effectiveness in improving regrasp planning efficiency. While there are other numerous publications, this paper does not focus heavily on this aspect and thus will not delve into further details.
@@
Another critical challenge in regrasp planning is uncertainty. Repeated regrasping can result in error accumulation, causing the grasp pose to deviate significantly from the planned outcome. The deviation can prevent the robot from successfully executing subsequent tasks. To address the uncertainty problems, Bauza et al. [21] integrated tactile information into regrasp planning and achieved precise and robust pick and placement for regrasping. Bronars et al. [22] proposed a simultaneous tactile estimation and control method to continuously plan sliding regrasp motion while maintaining grasping precisions. Gualtieri et al. [23] estimated the quality of grasp candidates against uncertainty based on point cloud completion, and optimized regrasp sequences to be robust. Zhang et al. [24] used reinforcement learning to learn a regrasp policy that optimizes grasp stability. Both vision and visual, tactile information were used for learning. This study can be partially classified as research aimed at reducing uncertainties in grasping poses during the regrasp planning process. However, our primary goal is not to achieve precise posse estimation for regrasping. Instead, we focus on actively leveraging the consecutive geometric constraints inherent in the regrasping process [25][26][27][28] to eliminate errors.
III Regrasp Planning Using Orthogonal Triples
The essence of using regrasp to eliminate uncertainty lies in leveraging the parallel finger pads of a gripper to constrain an object’s pose to a plane. By performing three successive regrasp using parallel-finger grippers along mutually orthogonal directions, an object can be constrained to three mutually orthogonal planes, and thus be precisely located. Therefore, the first step to achieve the goal of this study is to identify grasp poses along three mutually orthogonal directions. We are going to introduce the grasp planning and grouping methods in the first subsection below. Once the grasp poses and groups are determined, search and motion planning techniques can be employed to generate regrasp actions and consequently the robot motion required to eliminate uncertainties. The second subsection below details the specific methods for search and motion planning.
III-A Orthogonal Grasp Triples
@@
Subsequently, based on the results of the grasp planning, we proceed by separating them into groups. Since our goal is to identify three mutually orthogonal grasp poses, we set our grouping criterion to be based on the opening and closing directions of the gripper fingers. Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 1 Note 1 Note 1In this context, the direction of $\boldsymbol{v}_{i}$ is not considered. That is, $\boldsymbol{v}_{i}$ assumed identical to $-\boldsymbol{v}_{i}$.. Through the grouping process, we can obtain several groups represented by $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, …The candidate grasp poses in Fig. 2(b) can be right classified into three mutually orthogonal groups. They are visualized in Fig. 2(c) using red, green, and blue colors.
After obtaining the grasp groups, we perform triple combinations of the groups to identify three groups that satisfy the mutually orthogonal constraint. To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triple combination.
@@
The $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ in the above equation represent the opening and closing directions of three different groups in a triple. We sort the scores of all triples and prioritize them with a score closest to 0 for regrasp planning and uncertainty elimination. The closer the score value is to 00, the more orthogonal the three grasps from the groups are to each other. By performing successive regrasp using the prioritized triples, object uncertainties can be effectively minimized. The example shown in Fig. 2 represents a special case where the planned grasp poses are right divided into three orthogonal groups with a score of zero. In more general cases, there may be multiple candidate groups and triple combinations. The triples may have non-zero scores and regrasp should be performed by prioritizing the ones with smaller scores.
III-B Regrasp Planning
For the sorted triples, motion planning is sequentially performed on each of them, and the first successful solution is output as a valid candidate. Note that each triple comprises three groups of grasp poses. During planning, one pose from each group needs to be selected to form a combination of three grasps for computation. In this approach, we do not explicitly extract the three-grasp combinations as a separate step. Instead, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. The grasps in each group are sequentially assigned to the two robotic arms for planning. In this way, we incrementally explore the combinations of the grasp triples, and output successful motion once found. The detailed algorithmic flow is illustrated in Fig. 3. Note that the process illustrated in the figure is merely an incremental search. If a large amount of data can be accumulated using the proposed method in the future, we strongly recommend exploring learning-based methods to improve efficiency.
Figure 3: Regrasp planning used in this study is a simple incremental search method, which sequentially traverses grouped grasp triples until a valid solution is found. Since the regrasp planning algorithm itself is not the primary focus of this paper, further optimization aspects such as acceleration or grasp selection refinement are not explored.
@@ IV Conforming to Object Pose Uncertainty
Consider the scenario shown in Fig. 4(a), where the left hand has already moved the object from its initial pose to the handover pose, and the right hand is approaching to perform the regrasp. The target pose of the right hand is planned based on an assumed pose in the simulation. However, due to uncertainties in the real world, the object’s assumed pose in the simulation deviates from its actual pose, as illustrated by the orange dashed frames and red solid frames in Fig. 4(a). Consequently, the right hand’s grasping pose becomes misaligned with the actual object pose. Applying position control for grasping in this scenario may lead to undesirable outcomes. In mild cases, as shown in Fig. 4(b) and (c), the object may undergo deformation. In more severe cases, it may sustain damage. Therefore, rigidly executing the planned grasping motion solely with position control is not advisable.
@@
Figure 4: (a) The right hand approaches to grasp the object. Due to uncertainties, the object is not positioned exactly at the center of the two fingers as assumed but is instead slightly shifted upward. (b) At this point, if position control is used, the upper-side finger will make contact with the object first, exerting excessive force on it and causing deformation. (c) Comparison of object deformation before and after grasping.
@@
To address this issue, this study employs admittance control during robotic grasping and enables the manipulator’s pose to conform to the object’s pose. The basic concept of admittance control is to establish a mass-damper-spring system’s motion equation based on the external force exerted on the robot and the desired target force . The robot then follows the solution of this equation to adjust its motion accordingly, thereby accommodating external forces. For a desired robot position $x_{d}$, the admittance control equation is
$M(\ddot{\mathit{x}}_{p}-\ddot{\mathit{x}}_{d})+B(\dot{\mathit{x}}_{p}-\dot{\mathit{x}}_{d})+K(\mathit{x}_{p}-\mathit{x}_{d})=\mathit{f}+\mathit{f}_{d}.$ (2)
Here, $x_{p}$, $\dot{x_{p}}$, and $\ddot{x_{p}}$ are respectively the measured current robot end-effector position, velocity, and acceleration. $M$, $B$, $K$ denote the virtual inertia, damping, and stiffness parameters. $f$ is the force applied to the system, while $f_{d}$ represents the target force exerted by the robot after applying admittance control.
@@
In our specific grasping scenario, the target force is set to zero, as our objective is to ensure that the robot does not exert any external force on the object beyond the necessary grasping force, thereby preventing unintended movement, deformation, or damage. During the grasping process, when one of the robotic fingers first makes contact with the object, it receives a reaction force from the object, which corresponds to the external force $f$ in equation (2). Based on the robot’s current $x_{p}$, $\dot{x_{p}}$, $\ddot{x_{p}}$, as well as the force $f$ measured using a force/torque sensor installed at the robot manipulator’s wrist, we can compute the desired position $x_{d}$, velocity $\dot{x_{d}}$, and acceleration $\ddot{x_{d}}$ using the equation. These computed values are then used to control the robot, enabling it to conform to the object’s uncertain pose during the grasping process.
@@
Fig. 5 illustrates the effect of applying admittance control to the grasping process shown in Fig. 4. In Fig. 5(a), the upper-side finger first makes contact with the object and experiences a reaction force $f$. The admittance control then adjusts the robotic hand by moving it in a direction that balances this force, as shown in Fig. 5(b). The adjustment continues until both fingers are in contact with the object and their grasping forces reach equilibrium. At this point, $f$ disappears and the grasping forces become internal forces within the whole gripper-object system. The final grasping result is depicted in Fig. 4(c). Unlike position control, admittance control ensures that the object’s pose remains unchanged after grasping.
@@
Figure 5: Admittance control helps avoid deformation or damage. (a) When the upper-side finger makes contact with the object, a reaction force $f$ is detected. (b) Admittance control adjusts the robot hand by moving it in a direction that balances $f$. (c) After admittance control, the object is firmly grasped, and no deformation occurs in its shape.

### v3

# Bimanual Regrasp Planning and Control for Active Reduction of Object Pose Uncertainty

@@ Abstract
Precisely grasping an object is a challenging task due to pose uncertainties. Conventional methods have used cameras and fixtures to reduce object uncertainty. They are effective but require intensive preparation, such as designing jigs based on the object geometry and calibrating cameras with high-precision tools fabricated using lasers. In this study, we propose a method to reduce the uncertainty of the position and orientation of a grasped object without using a fixture or a camera. Our method is based on the concept that the flat finger pads of a parallel gripper can reduce uncertainty along its opening/closing direction through flat surface contact. Three orthogonal grasps by parallel grippers with flat finger pads collectively constrain an object’s position and orientation to a unique state. Guided by the concepts, we develop a regrasp planning and admittance control approach that sequentially finds and leverages three orthogonal grasps of two robotic arms to actively reduce uncertainties in the object pose. We evaluated the proposed method on different initial object uncertainties and verified that it had good repeatability. The deviation levels of the experimental trials were on the same order of magnitude as those of an optical tracking system, demonstrating strong relative inference performance.
@@
Grasping uncertainty, Regrasping, Bimanual manipulation
@@ I Introduction
Asignificant challenge in robotic manipulation lies in addressing the uncertainties associated with object grasping. The uncertainties often arise from errors in environmental registration, inaccuracies in object pose recognition, and unbalanced contact during grasping that leads to pose deviations. The uncertainties can result in discrepancies between the actual and expected poses of objects or tools, potentially causing task failures. Existing methods used to address uncertainty problems typically fall into two categories: Sensor-based correction methods [1][2] and fixture-based constraint methods [3][4]. Sensor-based methods leverage additional hardware, such as force sensors or tactile sensors, to detect and mitigate uncertainties. Fixture-based methods, on the other hand, rely on structuring the environment or using special-purpose structures to constrain object pose. However, both approaches have notable limitations, as they often require additional hardware adjustments or environment modifications.
Figure 1: (a) Precisely grasping an object is difficult. (a.i) Before grasping. (a.ii) After grasping. (a.iii) Comparison. (b) Flat finger pads of a parallel gripper help constrain uncertainty to a plane. (c) Three orthogonal parallel grasps constrain three planes, leading to an exact pose.
@@
This paper proposes a bimanual regrasp planning method that leverages flat finger pad constraints and admittance control to reduce uncertainties. It does not require external fixtures or additional sensors and is more flexible for diverse robotic manipulation scenarios. The fundamental concept of the proposed method is illustrated in Fig. 1(b) and (c). It is based on the observation that the flat finger pads of parallel grippers can reduce uncertainty along the hand’s opening/closing direction through flat surface contact. Three such grasps collectively constrain the object’s position and orientation to a unique state Note 1 Note 1 Note 1This unique pose constraint holds under the following assumptions: i) The object is sufficiently rigid such that the grasping force is transmitted and sensed through the force sensors, rather than causing deformation. ii) The contact between the gripper fingers and the object is assumed to be flat surface contact with sufficient friction to prevent slippage. iii) The maximum uncertainty in the object pose is within the gripper’s opening range, ensuring valid contact during grasping.. Guided by the concepts, we develop a regrasp planning approach [5] that sequentially executes a series of regrasping motions using two robotic arms while selecting mutually orthogonal grasp triplets to simultaneously reduce uncertainties in the object pose. However, grasp planning alone does not fully achieve our goal, as simply executing the planned motion sequence using position control may exert excessive grasping forces on the uncertain object and lead to deformation or even damage. To address this issue, we develop an admittance control [6] method that utilizes force/torque sensors embedded in the robot arm’s wrists to adjust the grasping pose and conform to the uncertain object pose. By integrating regrasp planning with admittance control, our method iteratively reduces uncertainties along three orthogonal axes, thus achieving an accurate object pose estimation.
In the experiments, we implemented the proposed method and evaluated its performance on different initial object uncertainties. The results demonstrate that the method achieves satisfactory repeatability. The deviation levels of remaining uncertainties were the same order of magnitude as those of an optical tracking system, demonstrating that the method provides strong relative inference performance. By leveraging the robot’s inherent geometric constraints and force sensors, the proposed approach can help actively reduce uncertainties in the grasped object.
@@ II Related Work
II-A Categories of Regrasp Planning Methods
Regrasp planning aims to automatically generate grasp sequences to accomplish object manipulation. It involves two main categories: discrete and continuous.
@@ II-A Categories of Regrasp Planning Methods
Discrete regrasp planning focuses on identifying discrete pick-and-place sequences. For example, Liu et al. [7] developed regrasp planning using multiple mobile robots. The method achieves flipping manipulation with the help of regrasping. Kim et al. [8] developed a randomized planning method for regrasp planning. Baek et al. [9] considered non-prehensile grasping in regrasp planning and secured power grasps for stable holding. Qin et al. [10] presented a regrasp planning method for installing a deformable linear object. Rigidity constraints were considered when planning regrasp motion. Mitrano et al. [11] studied a similar problem and proposed considering avoiding environment-grasp-body loops to accelerate grasp selection and planning. Sundaralingam et al. [12] extended regrasp planning to a multi-fingered hand. Murooka et al. [13] proposed the foot-step and regrasping planning method for a humanoid robot to simultaneously determine foot steps and hand switches.
Continuous regrasp methods, on the other hand, aim to adjust object poses through controlled rolling or sliding motions while maintaining contact. For instance, Maximo et al. [14] presented the grasp space method for continuous regrasp (rolling or sliding) that maintains force closures. Hu et al. [15] formulated regrasp planning as a multi-modal problem, where three manifolds, including the regrasp manifold, the transfer manifold, and the slide manifold, were planned across to determine action sequences and thus plan motion. Cheng et al. [16] further proposed the contact mode method to guide manipulation planning across multiple manifolds. The work presented in this paper is based on iterative pick-and-place actions using two robot hands, and thus it should be categorized under the first type of regrasp planning.
II-B Key Challenges and the Focus of this Work
@@
One of the most critical challenges in regrasp planning is the planning speed. Researchers have explored various methods to enhance efficiency. For example, Cho et al. [17] developed a fast regrasp planning algorithm using look-up tables and two-finger parallel grippers. Chiu et al. [18] proposed a reinforcement learning method to accelerate the planning of regrasp poses. Their actions were defined in the local frame of a robot hand, thereby resolving the dependencies on robotic kinematics. Wada et al. [19] significantly improved planning speed by using a neural network to filter waypoint poses and quickly select grasp poses. Xu et al. [20] proposed learning reorientation object poses and then planning regrasp motion while considering the learned middle reorientation poses. Learning-based methods have demonstrated effectiveness in improving regrasp planning efficiency. While there are numerous other publications, this paper does not focus heavily on this aspect and thus will not delve into further details.
@@
Another critical challenge in regrasp planning is uncertainty. Repeated regrasping can result in error accumulation, causing the grasp pose to deviate significantly from the planned outcome. The deviation can prevent the robot from successfully executing subsequent tasks. To address the uncertainty problems, Bauza et al. [21] integrated tactile information into regrasp planning and achieved precise and robust pick and placement for regrasping. Bronars et al. [22] proposed a simultaneous tactile estimation and control method to plan sliding regrasp motion while maintaining grasping precision continuously. Gualtieri et al. [23] estimated the quality of grasp candidates against uncertainty based on point cloud completion and optimized regrasp sequences to be robust. Zhang et al. [24] used reinforcement learning to learn a regrasp policy that optimizes grasp stability. Both vision and visual-tactile information were used for learning. This study can be partially classified as research to reduce uncertainties in grasping poses during the regrasp planning process. However, our primary goal is not to achieve precise pose estimation for regrasping. Instead, we focus on actively leveraging the consecutive geometric constraints inherent in the regrasping process [25][26][27][28] to reduce uncertainty.
From a broader perspective, our research also falls within the category of methods that utilize interactive perception [29] to estimate the pose of grasped objects. In addition to the grasping-based interaction approaches, recent literature on non-prehensile interaction has been receiving growing attention. For example, Jankowski et al. [30] discussed the stochastic contact dynamics of pushing and proposed a pushing planning method that minimizes uncertainty considering informed priors. Dong et al. [31], on the other hand, modeled the contact dynamics of pushing as an energy model and used constraints on the energy margin to plan robust pushing.
III Regrasp Planning Using Orthogonal Triplets
@@
Subsequently, based on the results of the grasp planning, we separate the grasps into groups. Since our goal is to identify three mutually orthogonal grasp poses, the grouping criterion is defined by the opening and closing directions of the gripper fingers. Grasp poses with the same opening and closing direction are grouped together, and each group can be represented by a unique opening and closing vector, denoted as $\boldsymbol{v}_{i}$ Note 2 Note 2 Note 2In this context, the direction of $\boldsymbol{v}_{i}$ is not considered.. Through the grouping process, we can obtain several groups represented by $\boldsymbol{v}_{1}$, $\boldsymbol{v}_{2}$, …. The candidate grasp poses in Fig. 2(b) can be correctly classified into three mutually orthogonal groups. They are visualized in Fig. 2(c) using red, green, and blue colors.
After obtaining the grasp groups, we perform triplet combinations of the groups to identify three groups that satisfy the mutually orthogonal constraint. To evaluate whether the constraint is satisfied, we use the following formula to compute a score for each triplet combination.
@@
The $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ in the above equation represent the opening and closing directions of three different groups in a triplet. We sort the scores of all triplets and prioritize them with a score closest to 0 for regrasp planning and uncertainty reduction. The closer the score value is to 00, the more orthogonal the three grasps from the groups are to each other. By performing successive regrasp using the prioritized triplets, object uncertainties can be effectively minimized.
Although we consistently refer to “orthogonal” grasps or groups, it is important to note that strict 90-degree orthogonality is not required to determine the object’s pose. The example shown in Fig. 2 represents a special case where the planned grasp poses are right divided into three orthogonal groups with a score of zero. In more general cases, there may be multiple candidate groups and triplet combinations. The triplets may have non-zero scores, However, as long as the grasp direction vectors $\boldsymbol{v}_{i}$, $\boldsymbol{v}_{j}$, and $\boldsymbol{v}_{k}$ form a non-singular matrix when arranged as columns, the pose can be estimated. This point will be clearer in Section V when we introduce the pose estimation algorithm. Here, the reason we sort triplet grasp combinations by their score is not for enforcing strict orthogonality, but rather to account for potential changes of grasp directions caused by subsequent admittance control. Prioritizing grasp triplets with higher orthogonality helps minimize the impact of grasp direction deviations. We will discuss this aspect in more detail in the admittance control section.
After identifying orthogonal groups, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. We use an incremental method to explore combinations of grasp triplets from the orthogonal groups and output a successful motion once found. Each triplet comprises three groups of grasp poses. During planning, one pose from each group needs to be selected to form a combination of three grasps for computation. In this approach, we do not explicitly extract the three-grasp combinations as a separate step. Instead, we leverage the two arms of a bimanual robot to perform regrasp planning interactively for the three grasp groups. The grasps in each group are sequentially assigned to the two robotic arms for planning. In this way, we incrementally explore the combinations of the grasp triplets and output successful motion once found. The detailed algorithmic flow is illustrated in Fig. 3.
Figure 3: Regrasp planning used in this study is a simple incremental search method, which sequentially traverses grouped grasp triplets until a valid solution is found.
It should be noted that our grasping and planning methods are probabilistic, and their success rates inherently depend on the granularity, number, and diversity of the grasp samples. At the same time, increasing the number of grasps inevitably impacts planning time. Identifying the optimal trade-off between these factors is an important challenge that could be addressed using predictive approaches. If a large amount of data can be accumulated using the proposed method in the future, we are interested in exploring probabilistic modeling methods to learn grasp triplets and thus improve planning efficiency. However, this is beyond the scope of the present paper and is therefore not elaborated here.
@@ IV Conforming to Object Pose Uncertainty
Consider the scenario shown in Fig. 4(a), where the giving hand has already moved the object from its initial pose to the handover pose, and the receiving hand is approaching to perform the regrasp. The target pose of the giving hand is planned based on an assumed pose in the simulation. However, due to uncertainties in the real world, the object’s assumed pose in the simulation deviates from its actual pose, as illustrated by the orange dashed frames and red solid frames in Fig. 4(a). Consequently, the receiving hand’s grasping pose becomes misaligned with the actual object pose. Applying position control for grasping in this scenario may lead to undesirable outcomes. In mild cases, as shown in Fig. 4(b) and (c), the object may undergo deformation. In more severe cases, it may sustain damage. Therefore, rigidly executing the planned grasping motion solely with position control is not advisable.
@@
Figure 4: (a) The receiving hand approaches to grasp the object. Due to uncertainties, the object is not positioned exactly at the center of the two fingers, as assumed, but is instead slightly shifted upward. (b) At this point, if position control is used, the upper-side finger will make contact with the object first, exerting excessive force on it and causing deformation. (c) Comparison of object deformation before and after grasping.
@@
To address this issue, this study employs admittance control during robotic grasping and enables the receiving hand’s grasping pose to conform to the uncertain object pose. Admittance control establishes a mass-damper-spring system’s motion equation based on the reaction force exerted on the receiving hand and the desired target force [6]. The receiving hand follows the solution of the equation to adjust its motion accordingly, thereby accommodating reaction force. For a desired receiving hand position $\boldsymbol{x}_{d}$, the admittance control equation is
$\boldsymbol{M}(\ddot{\boldsymbol{x}}_{p}-\ddot{\boldsymbol{x}}_{d})+B(\dot{\boldsymbol{x}}_{p}-\dot{\boldsymbol{x}}_{d})+K(\boldsymbol{x}_{p}-\boldsymbol{x}_{d})=\boldsymbol{f}+\boldsymbol{f}_{d}.$ (2)
@@
Here, $\boldsymbol{x}_{p}$, $\dot{\boldsymbol{x}_{p}}$, and $\ddot{\boldsymbol{x}_{p}}$ are respectively the measured current end-effector position, velocity, and acceleration of the receiving hand. $\boldsymbol{M}$, $\boldsymbol{B}$, and $\boldsymbol{K}$ denote the virtual inertia, damping, and stiffness parameters. $\boldsymbol{f}$ is the reaction force applied to the receiving hand when grasping. It can be measured using the force/torque sensor at the wrist of the receiving manipulator. $\boldsymbol{f}_{d}$ represents the target exerting force of the receiving hand after applying admittance control. It is essentially zero, as our goal is to ensure that the receiving hand does not exert any external force on the object beyond the necessary grasping force, thereby preventing unintended movement, deformation, or damage. Given the robot’s current $\boldsymbol{x}_{p}$, $\dot{\boldsymbol{x}_{p}}$, $\ddot{\boldsymbol{x}_{p}}$, $\boldsymbol{f}$, the desired trajectory $\boldsymbol{x}_{d}$, $\dot{\boldsymbol{x}}_{d}$, and $\ddot{\boldsymbol{x}}_{d}$ can be computed from the admittance control equation. These computed values are then fed to the admittance control law to drive the receiving hand to conform to the object’s uncertain pose during grasping.
@@
Fig. 5 illustrates the effect of applying admittance control to the grasping process shown in Fig. 4. In Fig. 5(a), the upper-side finger first makes contact with the object and experiences a reaction force $\boldsymbol{f}$. The admittance control then adjusts the robotic hand by moving it in a direction that balances this force, as shown in Fig. 5(b). The adjustment continues until both fingers are in contact with the object and their grasping forces reach equilibrium. At this point, $\boldsymbol{f}$ disappears, and the grasping forces become internal forces within the whole gripper-object system. The final grasping result is depicted in Fig. 4(c). Unlike position control, admittance control ensures that the object’s pose remains unchanged after grasping.
@@
Figure 5: Admittance control helps avoid deformation or damage. (a) A reaction force $\boldsymbol{f}$ is detected when the upper-side finger contacts the object. (b) Admittance control adjusts the receiving hand by moving it to balance $\boldsymbol{f}$. (c) After admittance control, the object is firmly grasped without change.
Note that the above equation describes only the position-based admittance control. A similar formulation also applies to rotational admittance control. However, the admittance controller may redirect the grasp toward an adjacent surface due to rotational errors. In extreme cases, this can result in parallel regrasps, making the object’s pose indeterminate. As a safeguard against such situations, we suggest prioritizing near-orthogonal grasps following equation (1). The design minimizes the likelihood that the three grasps lie on the same plane, even when one of the grasp directions is redirected due to unexpected admittance behavior.